# Graph Max Shift: A Hill-Climbing Method for Graph Clustering

### v1

@@ 6.2 Auxiliary results
Since we are assuming that $\eta<s_{2}-s_{1}$, it is the case that $s_{2}-\eta>s_{1}$, so that $\mathcal{C}_{s_{2}-\eta}(x^{*})\subset\mathcal{C}_{s_{1}}(x^{*})$. Assume that $\epsilon$ is strictly smaller than the minimum separation between the connected components of $\mathcal{U}_{s_{1}}$ and consider Graph Max Shift initialized at some $y_{0}\in B\left(x^{*},\delta_{2}\right)$. Note that $y_{0}\in\hat{\mathcal{C}}_{s_{2}}(x^{*})$, and by the fact that $(y_{k})$ is hill-climbing with respect to $\hat{f}$, it must be the case that $(y_{k})\subset\hat{\mathcal{U}}_{s_{2}}$. In fact, since by Lemma 6.2 each connected component of $\hat{\mathcal{U}}_{s_{2}}$ must be within a connected component of $\mathcal{U}_{s_{2}-\eta}$, and since $s_{2}-\eta>s_{1}$, each connected component of $\mathcal{U}_{s_{2}-\eta}$ must be within a connected component of $\mathcal{U}_{s_{1}}$, and these are separated by a distance exceeding $\epsilon$, and since the increments in $(y_{k})$ are of length bounded from above by $\epsilon$ by Lemma 6.6, it must be the case that $(y_{k})\subset\hat{\mathcal{C}}_{s_{2}}(x^{*})$, that is, the sequence must remain in the same connected component of $\hat{\mathcal{U}}_{s_{2}}$. Continuing, since there are a finite number of points in $\mathcal{Y}$, the sequence $(y_{k})$ must converge in a finite number of steps. Let $y_{\infty}$ denote the limit, which by the fact that $(y_{k})\subset\hat{\mathcal{C}}_{s_{2}}(x^{*})$ and (6.35) above must be in $B(x^{*},\delta_{1})$. Therefore, due to (6.34),
$\|\nabla f(y_{\infty})\|\geq\lambda_{0}\|y_{\infty}-x^{*}\|.$ (6.36)
We now take $s_{1}$ large enough that $s_{1}\geq s^{*}/2$, and we let $\alpha$ be short for $\alpha_{s/2}$. Define $z=y_{\infty}+(\epsilon-\alpha)N(y_{\infty})$, and let $y$ be the closest point in $\mathcal{Y}$ to $z$. By (6.2) and the fact that $y_{\infty}\in\mathcal{C}_{s_{2}-\eta}$,
$f(z)\geq f(y_{\infty})-\kappa_{1}\epsilon\geq s_{2}-\eta-\kappa_{1}\epsilon,$ (6.37)
so that $f(z)\geq s_{1}$ if $\epsilon$ and $\eta$ are small enough that $\kappa_{1}\epsilon+\eta\leq s_{2}-s_{1}$, which we assumed henceforth. Because of that, $\|y-z\|\leq\alpha$, implying via the triangle inequality that $\|y-y_{\infty}\|\leq\epsilon$, making $y$ a fair search point when the sequence is at $y_{\infty}$, and for that reason, it must be the case that $\hat{f}(y_{\infty})\geq\hat{f}(y)$. Assuming that $\epsilon>\alpha$, we then derive
$0\geq\hat{f}(y)-\hat{f}(y_{\infty})$ $\geq f(y)-f(y_{\infty})-2\eta$ (6.38)
$\geq f(z)-f(y_{\infty})-\kappa_{1}\alpha-2\eta$ (6.39)
$\geq\nabla f(y_{\infty})^{\top}(z-y_{\infty})-\kappa_{2}\|z-y_{\infty}\|^{2}-\kappa_{1}\alpha-2\eta$ (6.40)
$=\|\nabla f(y_{\infty})\|(\epsilon-\alpha)-\kappa_{2}(\epsilon-\alpha)^{2}-\kappa_{1}\alpha-2\eta$ (6.41)
$\geq\|\nabla f(y_{\infty})\|\epsilon-\kappa_{2}\epsilon^{2}-2\alpha\kappa_{1}-2\eta.$ (6.42)
@@
$\|y_{\infty}-x^{*}\|\leq\frac{\kappa_{2}\epsilon^{2}+2\alpha\kappa_{1}+2\eta}{\lambda_{0}\epsilon},$
which establishes the claim. ∎
$x_{k+1}\in\operatorname*{arg\,max}\limits_{x\in\bar{B}(x_{k},\epsilon)}f(x),\quad k\geq 0.$ (6.43)
The first two results are variants of Lemma 6.6 and Lemma 6.7, respectively, with $f$ in place of $\hat{f}$ and a search not restricted to $\mathcal{Y}$.
Lemma 6.9.
For the Max Shift sequence $(x_{k})$ initialized at any $x_{0}$,
$\epsilon-C_{0}\frac{\epsilon^{2}}{\|\nabla f(x_{k})\|}\leq\|x_{k+1}-x_{k}\|\leq\epsilon,\quad\forall k\geq 0.$ (6.44)
Lemma 6.10.
For the Max Shift sequence $(x_{k})$ initialized at any $x_{0}$,
$x_{k+1}-x_{k}=\|x_{k+1}-x_{k}\|N(x_{k})\pm 2C_{0}\frac{\epsilon^{3/2}}{\|\nabla f(x_{k})\|^{1/2}},\quad\forall k\geq 0.$ (6.45)
The third result is distilled from the proof of Th 3.1–3.2 and arguments provided in Sec 3.2.1 in the same paper [2].
Lemma 6.11.
For any $p>0$, there is $s>0$ and a subset $\Omega\subset\mathcal{U}_{s}$ with $\int_{\Omega}f\geq 1-p$ such that any point in $\Omega$ is in the basin of attraction of a mode. Moreover, for any $\delta>0$, there is $C_{2}>0$ such that, if $\epsilon\leq 1/C_{2}$, starting at any point $x_{0}\in\Omega$ in the basin of attraction of some mode $x^{*}$, the sequence $(x_{k})$ that Max Shift computes enters $B(x^{*},\delta)$ in at most $C_{2}/\epsilon$ steps, while $\|\nabla f(x_{k})\|\geq 1/C_{2}$ for all $k\geq 0$ such that $x_{k}\notin B(x^{*},\delta)$.
6.3 Proof of Theorem 3.3
It is enough to show that, for any $p>0$, there is $s>0$ and a subset $\Omega\subset\mathcal{U}_{s}$ with $\int_{\Omega}f\geq 1-p$ such that, with probability at least $1-p$ under the conditions of the theorem, any data point in $\Omega$ that is in the basin of attraction of a mode is moved by Graph Max Shift to a point within distance $\frac{1}{3}\tau\epsilon$ of that mode. For that to work, we assume that $\alpha_{s/2}/\epsilon$ is sufficiently small that $\alpha_{s/2}/\epsilon<1/2C$ where $C$ is the constant of Lemma 6.5, which is implied by our assumptions (see further down). In that case, with probability at least $1-p$, if we take two data points in $\Omega$ (representing an expected proportion of $(1-p)^{2}$ of all pairs of data points), if they are in the basin of attraction of the same mode, they will be moved by Graph Max Shift to points within distance $\frac{2}{3}\tau\epsilon$, so that the graph distance between the corresponding nodes is bounded by $(1+C\alpha_{s/2}/\epsilon)\frac{2}{3}\tau\epsilon+1<\tau+1$ by Lemma 6.5, implying a number of hops $\leq\tau$, and therefore clustered together; and if they are in different basins of attraction, assuming that $\tau\epsilon$ is smaller than half the minimum separation between any two modes in $\mathcal{U}_{s}$, they will be moved by Graph Max Shift to points within distance $\frac{1}{3}\tau\epsilon$ of their associated modes, so that their distance will exceed $2\tau\epsilon-\frac{2}{3}\tau\epsilon>\tau\epsilon$, implying that the corresponding clusters will not be merged. (Note that, regardless of $s>0$, there are only finitely many modes within $\mathcal{U}_{s}$.)
Therefore, fix $p>0$, let $s>0$ and $\Omega\subset\mathcal{U}_{s}$ be as in Lemma 6.11. Let $\alpha$ be short for $\alpha_{s/2}$. For $\epsilon\leq q\leq 1$ a constant to be chosen small later on, consider the event
$q\epsilon\geq\max\big{\{}\alpha,\eta\big{\}}.$ (6.46)
This event happens with probability tending to 1 because of the last condition on $\epsilon$ in the statement of the theorem, as well as Lemma 6.3 and Lemma 6.4, which give the order of magnitude of $\eta$ and $\alpha$. We assume that $\epsilon$ is small enough that all the lemmas in Section 6.2, called upon below, apply and that (6.46) occurs with probability at least $1-p$.
Note that we have already imposed a requirement on $q$ above, that $q<1/2C$, where $C$ is the constant of Lemma 6.5, so that $\alpha/\epsilon<1/2C$. Below, each time there is a constraint on $q$, we assume it is small enough to satisfy the constraint. We just have to make sure that each constraint allows for $q>0$ fixed and that there are only a finite (in fact, a few) constraints.
Taking any data point $y_{0}\in\mathcal{Y}\cap\Omega$ as a starting point, let $(y_{k})$ denote the sequence computed by Graph Max Shift based on $\hat{f}$ as defined in (6.15); let $(x_{k})$ denote the sequence computed by Max Shift based on $f$ as defined in (6.43) (with $x_{0}=y_{0}$). We note that, combining (6.16) and (6.17), and under (6.46), we have
$y_{k+1}-y_{k}$ $=\epsilon N(y_{k})\pm C_{0}\frac{\epsilon^{2}+\alpha+\eta}{\|\nabla f(y_{k})\|}\pm 2C_{0}\frac{\epsilon^{1/2}(\epsilon^{2}+\alpha+\eta)^{1/2}}{\|\nabla f(y_{k})\|^{1/2}}$ (6.47)
$=\epsilon N(y_{k})\pm C^{\prime}_{0}\frac{q^{1/2}\epsilon}{\|\nabla f(y_{k})\|},$ (6.48)
using the fact that $\|\nabla f(y_{k})\|\leq\kappa_{1}$. Similarly, combining (6.44) and (6.45), we have
$x_{k+1}-x_{k}$ $=\epsilon N(x_{k})\pm C^{\prime}_{0}\frac{q^{1/2}\epsilon}{\|\nabla f(x_{k})\|}.$ (6.49)
The inequalities above are valid for all $k\geq 0$.
Let $x^{*}$ be the mode to which $y_{0}$ is attracted by gradient ascent. Let $\delta$ be as in Lemma 6.8, and suppose we are able to show that
$y_{k}\in\bar{B}(x^{*},\delta)$ for some $k$. (6.50)
Then Lemma 6.8 tells us that $(y_{k})$ will end at some point within distance $C_{1}(\epsilon+(\alpha+\eta)/\epsilon)$ of $x^{*}$. Based on our assumptions on $\tau$ and $\epsilon$ in the statement of the theorem, this is bounded by $\frac{1}{3}\tau\epsilon$ eventually in the limit $n\to\infty$. Therefore, it suffices to show that (6.50) happens under (6.46) when $q$ is small enough. Assuming that $x_{0}=y_{0}\notin B(x^{*},\delta/2)$, for otherwise our job is done, let $k_{\#}$ be such that $x_{k_{\#}}\in B(x^{*},\delta/2)$ while $x_{k}\notin B(x^{*},\delta/2)$ for $k<k_{\#}$, which is well-defined by Lemma 6.11, which also tells us that $k_{\#}\leq C_{2}/\epsilon$. By the triangle inequality, to establish (6.50) it is enough to show that
$\|x_{k_{\#}}-y_{k_{\#}}\|\leq\delta/2.$ (6.51)
Define $C^{\prime}_{1}=2C_{2}C^{\prime}_{0}$ and $C^{\prime}_{2}=(2C_{2}+1)\kappa_{2}$, and assume $q$ is small enough that
$q^{1/2}\big{[}2C^{\prime}_{1}{C^{\prime}_{2}}^{-1}(\exp(C^{\prime}_{2}C_{2})-1)\big{]}\leq\frac{1}{2C_{2}\kappa_{2}}.$
We show by recursion that
$\|x_{k}-y_{k}\|\leq q^{1/2}[2C^{\prime}_{1}{C^{\prime}_{2}}^{-1}](\exp(C^{\prime}_{2}\epsilon k)-1),$ (6.52)
for all $k\leq k_{\#}$. By the fact that $k_{\#}\leq C_{2}/\epsilon$, when the inequality (6.52) is satisfied, we also have
$\|x_{k}-y_{k}\|\leq\frac{1}{2C_{2}\kappa_{2}}.$ (6.53)
We start by verifying that (6.52) is true at $k=0$ since $x_{0}=y_{0}$. Assume it is true at some $k<k_{\#}$, so that (6.53) is also true, and by Lemma 6.11 and a Taylor development,
$\|\nabla f(y_{k})\|\geq\|\nabla f(x_{k})\|-\kappa_{2}\|x_{k}-y_{k}\|\geq\frac{1}{C_{2}}-\kappa_{2}\frac{1}{2C_{2}\kappa_{2}}=\frac{1}{2C_{2}}.$
Combining this with (6.48) and (6.49), we find that
$y_{k+1}-y_{k}$ $=\epsilon N(y_{k})\pm C^{\prime}_{1}q^{1/2}\epsilon;$ (6.54)
$x_{k+1}-x_{k}$ $=\epsilon N(x_{k})\pm C^{\prime}_{1}q^{1/2}\epsilon.$ (6.55)
We then have, by the triangle inequality,
$\|x_{k+1}-y_{k+1}\|$ $\leq\|x_{k}-y_{k}\|+\epsilon\|N(x_{k})-N(y_{k})\|+2C^{\prime}_{1}q^{1/2}\epsilon.$ (6.56)
In view of (6.6), and the fact (justified by the same Taylor development) that $\|\nabla f(z)\|\geq 1/2C_{2}$ for all $z$ on the line segment joining $x_{k}$ and $y_{k}$,
$\|N(x_{k})-N(y_{k})\|\leq C^{\prime}_{2}\|x_{k}-y_{k}\|,$
implying that
$\|x_{k+1}-y_{k+1}\|$ $\leq(1+C^{\prime}_{2}\epsilon)\|x_{k}-y_{k}\|+2C^{\prime}_{1}q^{1/2}\epsilon.$ (6.57)
Using the recursion hypothesis (6.52), we find after some simplifications based on $e^{a}-1-a\geq 0$ for all $a$, we deduce
$\|x_{k+1}-y_{k+1}\|\leq q^{1/2}[2C^{\prime}_{1}{C^{\prime}_{2}}^{-1}](\exp(C^{\prime}_{2}\epsilon(k+1))-1),$ (6.58)
so that the recursion carries on.
In particular, we have established that
$\|x_{k_{\#}}-y_{k_{\#}}\|\leq q^{1/2}[2C^{\prime}_{1}{C^{\prime}_{2}}^{-1}](\exp(C^{\prime}_{2}C_{2})-1),$ (6.59)
and taking $q$ small enough that the term on the right-hand side is bounded by $\delta/2$ results in $\|x_{k_{\#}}-y_{k_{\#}}\|\leq\delta/2$, which is what we stated in (6.51) that we needed to establish. ∎

### v2

@@ 6.2 Results about Max Shift
Since we are assuming that $\eta<s_{2}-s_{1}$, it is the case that $s_{2}-\eta>s_{1}$, so that $\mathcal{C}_{s_{2}-\eta}(x^{*})\subset\mathcal{C}_{s_{1}}(x^{*})$. Assume that $\epsilon$ is strictly smaller than the minimum separation between the connected components of $\mathcal{U}_{s_{1}}$ and consider Graph Max Shift initialized at some $y^{0}\in B\left(x^{*},\delta_{2}\right)$. Note that $y^{0}\in\hat{\mathcal{C}}_{s_{2}}(x^{*})$, and by the fact that $(y^{k})$ is hill-climbing with respect to $\hat{f}$, it must be the case that $(y^{k})\subset\hat{\mathcal{U}}_{s_{2}}$. In fact, since by Lemma 6.2 each connected component of $\hat{\mathcal{U}}_{s_{2}}$ must be within a connected component of $\mathcal{U}_{s_{2}-\eta}$, and since $s_{2}-\eta>s_{1}$, each connected component of $\mathcal{U}_{s_{2}-\eta}$ must be within a connected component of $\mathcal{U}_{s_{1}}$, and these are separated by a distance exceeding $\epsilon$, and since the increments in $(y^{k})$ are of length bounded from above by $\epsilon$, it must be the case that $(y^{k})\subset\hat{\mathcal{C}}_{s_{2}}(x^{*})$, that is, the sequence must remain in the same connected component of $\hat{\mathcal{U}}_{s_{2}}$. Continuing, since there are a finite number of points in $\mathcal{Y}$, the sequence $(y^{k})$ must converge in a finite number of steps. Let $y^{\infty}$ denote the limit, which by the fact that $(y^{k})\subset\hat{\mathcal{C}}_{s_{2}}(x^{*})$ and (6.56) above must be in $B(x^{*},\delta_{1})$. Therefore, due to (6.55),
$\|\nabla f(y^{\infty})\|\geq\lambda_{0}\|y^{\infty}-x^{*}\|.$ (6.57)
We now take $s_{1}$ large enough that $s_{1}\geq s^{*}/2$, and we let $\alpha$ be short for $\alpha_{s/2}$. Define $z=y^{\infty}+(\epsilon-\alpha)N(y^{\infty})$, and let $y$ be the closest point in $\mathcal{Y}$ to $z$. By (6.2) and the fact that $y^{\infty}\in\mathcal{C}_{s_{2}-\eta}$,
$f(z)\geq f(y^{\infty})-\kappa_{1}\epsilon\geq s_{2}-\eta-\kappa_{1}\epsilon,$ (6.58)
so that $f(z)\geq s_{1}$ if $\epsilon$ and $\eta$ are small enough that $\kappa_{1}\epsilon+\eta\leq s_{2}-s_{1}$, which we assume henceforth. Because of that, $\|y-z\|\leq\alpha$, implying via the triangle inequality that $\|y-y^{\infty}\|\leq\epsilon$, making $y$ a fair search point when the sequence is at $y^{\infty}$, and for that reason, it must be the case that $\hat{f}(y^{\infty})\geq\hat{f}(y)$. It is enough to prove the result when $\epsilon>\alpha$, and assuming this is the case, we derive
$0\geq\hat{f}(y)-\hat{f}(y^{\infty})$ $\geq f(y)-f(y^{\infty})-2\eta$ (6.59)
$\geq f(z)-f(y^{\infty})-\kappa_{1}\alpha-2\eta$ (6.60)
$\geq\nabla f(y^{\infty})^{\top}(z-y^{\infty})-\kappa_{2}\|z-y^{\infty}\|^{2}-\kappa_{1}\alpha-2\eta$ (6.61)
$=\|\nabla f(y^{\infty})\|(\epsilon-\alpha)-\kappa_{2}(\epsilon-\alpha)^{2}-\kappa_{1}\alpha-2\eta$ (6.62)
$\geq\|\nabla f(y^{\infty})\|\epsilon-\kappa_{2}\epsilon^{2}-2\alpha\kappa_{1}-2\eta.$ (6.63)
@@
$\|y^{\infty}-x^{*}\|\leq\frac{\kappa_{2}\epsilon^{2}+2\alpha\kappa_{1}+2\eta}{\lambda_{0}\epsilon},$
and from this we conclude. ∎
6.3.1 Proof of (i)
Let $\rho$ be the quantity on the LHS of (3.6). Given $p>0$, let $s,\Omega,A$ be as in Proposition 6.9, and consider the quantity
$\rho_{\Omega}:=\frac{\#\big\{i<j:\widehat{\mathbf{I}}(i,j)<\mathbf{I}(i,j);\ y_{i},y_{j}\in\Omega\big\}}{\binom{n}{2}}$, (6.64)
where $\widehat{\mathbf{I}}$ is the output partition of Graph Max Shift without a merging post-processing step. By definition, $\rho_{\Omega}\leq\rho$, and with $\Omega:=\#\{i:y_{i}\in\Omega\}$, we also have
$\rho-\rho_{\Omega}$ $\leq\#\big\{i<j:y_{i}\notin\Omega\text{ or }y_{j}\notin\Omega\big\}\binom{n}{2}^{-1}$ (6.65)
$=1-\#\Omega(\#\Omega-1)/n(n-1)$ (6.66)
$\to 1-(\textstyle\int_{\Omega}f)^{2}$ (6.67)
$\leq 1-(1-p)^{2},$ (6.68)
in probability as $n\to\infty$, since $\Omega\sim\text{Bin}(n,\int_{\Omega}f)$ and $\int_{\Omega}f\geq 1-p$. By Lemma 6.3 and our assumption on $\epsilon$ in (i), we have $\eta=o_{P}(\epsilon)$; and by Lemma 6.4, letting $\alpha=\alpha_{s/2}$, we also have $\alpha=o_{P}(\epsilon)$. Therefore, $A(\epsilon+(\alpha+\eta)/\epsilon)=o_{P}(1)$. Let $\delta$ be the minimum separation between two modes within $\mathcal{U}_{s/2}$ — which is positive since $\mathcal{U}_{s/2}$ is compact and the modes are isolated — and let $\mathcal{E}$ be the event that $A(\epsilon+(\alpha+\eta)/\epsilon)<\delta/2$. Then $\operatorname{\mathbb{P}}(\mathcal{E})\to 1$ as $n\to\infty$, and under $\mathcal{E}$, $\rho_{\Omega}=0$ by Proposition 6.9. Bringing all together, we have argued that $\limsup\rho\leq 1-(1-p)^{2}$ in probability as $n\to\infty$, and given that $p$ can be chosen arbitrarily close to zero, we conclude that $\rho\to 0$ in probability.
6.3.2 Proof of (ii)
The arguments are, at first, completely parallel with those in Section 6.3.1. We redefine the notation used there to accommodate the setting of (ii). Let $\rho$ be the quantity on the LHS of (3.7). Given $p>0$, let $s,\Omega,A$ be as in Proposition 6.9, and consider the quantity
$\rho_{\Omega}:=\frac{\#\big\{i<j:\widehat{\mathbf{I}}(i,j)\neq\mathbf{I}(i,j);\ y_{i},y_{j}\in\Omega\big\}}{\binom{n}{2}}$, (6.69)
where $\widehat{\mathbf{I}}$ is the output partition of Graph Max Shift with a merging post-processing step with parameter $\tau\geq 1$. By definition, $\rho_{\Omega}\leq\rho$, and as in Section 6.3.1,
$\rho-\rho_{\Omega}$ $\leq 1-\#\Omega(\#\Omega-1)/n(n-1)$ (6.70)
$\to 1-(\textstyle\int_{\Omega}f)^{2}$ (6.71)
$\leq 1-(1-p)^{2},$ (6.72)
in probability as $n\to\infty$. By Lemma 6.3 and our assumption on $\epsilon$ in (ii), we have $\eta=o_{P}(\epsilon^{2})$; and by Lemma 6.4, letting $\alpha=\alpha_{s/2}$, we also have $\alpha=o_{P}(\epsilon^{2})$. Therefore, $A(\epsilon+(\alpha+\eta)/\epsilon)=A(\epsilon+o_{P}(\epsilon))$. (Compare with Section 6.3.1.) Let $\mathcal{E}$ be the event that $A(\epsilon+(\alpha+\eta)/\epsilon)\leq 2A\epsilon$, so that $\operatorname{\mathbb{P}}(\mathcal{E})\to 1$ as $n\to\infty$. First, eventually, $2A\epsilon<\delta/2$, where $\delta$ is as in Section 6.3.1, and when this is the case, under $\mathcal{E}$, as in Section 6.3.1,
$\#\big\{i<j:\widehat{\mathbf{I}}(i,j)<\mathbf{I}(i,j);\ y_{i},y_{j}\in\Omega\big\}=0.$ (6.73)
Further, suppose that $\tau\geq 4A$. By Proposition 6.9, points that belong to the same basin of attraction within $\Omega$ are moved to terminal nodes within distance $2A\epsilon$ of the corresponding mode, and therefore within $4A\epsilon$ of each other by the triangle inequality, so that they are grouped together in post-processing, implying that we also have
$\#\big\{i<j:\widehat{\mathbf{I}}(i,j)>\mathbf{I}(i,j);\ y_{i},y_{j}\in\Omega\big\}=0.$ (6.74)
We thus have $\rho_{\Omega}=0$, and we may conclude as we did in Section 6.3.1.
Appendix A Additional Plots
We depict the weak clustering error (3.6) and clustering error (3.7) as a function of $\epsilon$ for the other Gaussian mixture models considered in Figure 4.1. Observe that the quadrimodal density for exhibits higher variance than the bimodal density, for example.
Figure A.1: The weak clustering error (left) and the clustering error (right) of the other Gaussian mixtures from Figure 4.1 when Graph Max Shift is applied to $\mathcal{G}(\mathcal{Y},\epsilon)$. The tuning parameter $\tau=1$ for all experiments. Solid lines indicate the error averaged over $100$ simulations, and the shaded regions indicate empirical 80% intervals.
We additionally include another bimodal Gaussian mixture. In this mixture, the separation between the two components is small so that the clustering task is more difficult. In Figure A.2, we plot the clustering produced by Graph Max Shift with $\tau=1$ over a range of values of $\epsilon$. Note that while the weak clustering error is small when $\epsilon=0.035$ and $\epsilon=0.06$, but the clusters are over-segmented in these cases. When $\epsilon$ is larger, both the weak clustering error and clustering error are large, as some points have crossed the basins of attraction. However, using the tuning parameter on the same sample, a small clustering error can be achieved. This is depicted in Figure A.3, where $\tau$ is adjusted with $\epsilon$ fixed. When $\tau$ is too small the left cluster is split into multiple clusters, but by increasing $\tau=2$, the nodes in the left component are merged together and most points are correctly clustered. When $\tau$ is too large, clusters in different basins of attraction are merged together.
Figure A.2: Graph Max Shift applied to $\mathcal{G}(\mathcal{Y};\epsilon)$ with data drawn from a bimodal Gaussian mixture. The top row shows the obtained clustering with the indicated $\epsilon$ and $\tau=1$. The bottom row depicts the degree, which is proportional to the value of the density estimator implicitly computed. Additionally, each plot includes a ball of radius $\epsilon$ in the bottom right for reference.
Figure A.3: Graph Max Shift applied to $\mathcal{G}(\mathcal{Y};\epsilon)$ with data drawn from a bimodal Gaussian mixture. The obtained clustering with $\epsilon=0.06$ fixed as $\tau$ is varied in $\{1,2,3\}$. The nodes associated with the clusters are plotted in black. Additionally, each plot includes a ball of radius $\epsilon$ in the bottom right for reference.